A buffer system (pKa = 5) is composed of 0.1 moles of weak acid and 0.05 moles of the conjugate base. What is the pH of the system after the addition of 0.07 moles of a strong base?
Answer #1

after adding strong acid following reaction takes place

strong base + weak acid -----------> conjugate base + water

after strong acid added

assume total volume = 1.0 L

moles of conjugate base = 0.05 + 0.07 = 0.12

moles of weak acid = 0.1 - 0.07 = 0.03

[weak acid] = 0.03 M

[conjugate base] = 0.12 M

for acidic buffer

pH = pKa + log [conjugate base] / [weak acid]

pH = 5.0 + log [0.12] / [0.03]

pH = 5.60